# Breaking the Gaussian Barrier: Residual-PAC Privacy for Automatic Privatization

### v1

@@ 3.1 Foundation of Residual-PAC Privacy
In this section, we develop general results to support concrete analyses under the Residual-PAC Privacy framework. We begin by introducing key information-theoretic quantities, entropy and conditional entropy, which are fundamental to our development.
Definition 6 (Entropy and Conditional Entropy).
Let $X$ be a random variable on a discrete alphabet $\mathcal{X}$ with probability mass function $P_{X}(x)=\Pr(X=x)$. The Shannon entropy of $X$ is
$\mathcal{H}(X)=-\sum\nolimits_{x\in\mathcal{X}}P_{X}(x)\,\log P_{X}(x).$
If $X$ instead has a continuous distribution on $\mathcal{X}\subseteq\mathbb{R}^{n}$ with probability density function $f_{X}(x)$, its differential entropy is
[…]
Moreover, let $(X,W)$ be jointly distributed. If $X,W$ are discrete with joint PMF $p_{X,W}(x,w)$ and marginals $p_{W}(w)$, the conditional entropy of $X$ given $W$ is
$\mathcal{H}(X|W)$ $=\sum\nolimits_{w\in\mathcal{W}}P_{W}(w)\,\Bigl{[}$
$-\sum\nolimits_{x\in\mathcal{X}}P_{X|W}(x\mid w)\,\log P_{X|W}(x|w)\Bigr{]}.$
If $(X,W)$ are continuous with joint PDF $f_{X,W}(x,w)$ and marginal $f_{W}(w)$, the differential conditional entropy is
$h(X|W)$ $=\int_{\mathcal{W}}f_{W}(w)\,\Bigl{[}$
$-\int_{\mathcal{X}}f_{X|W}(x|w)\,\log f_{X\mid W}(x|w)\,dx\Bigr{]}\,dw.$
For ease of exposition, we use $\mathcal{H}(X)$ to denote the entropy of $X$, either Shannon or differential depending on the context, and $\mathcal{H}(X|W)$ to denote the corresponding conditional entropy.
Mutual information plays a central role in the PAC Privacy analysis, as it is defined in terms of entropy and conditional entropy. Let $X$ and $W$ be random variables with joint distribution $P_{X,W}(x,w)$ and marginals $P_{X}(x)$ and $P_{W}(w)$. Then, the mutual information between $X$ and $W$ is given by
$\mathtt{MI}(X;W)=\mathcal{H}(X)-\mathcal{H}(X|W)=\mathtt{D}_{\mathrm{KL}}\bigl{(}P_{X,W}\,\|\,P_{X}\otimes P_{W}\bigr{)},$
where $\mathtt{D}_{\mathrm{KL}}\bigl{(}P_{X,W}\,\|\,P_{X}\otimes P_{W}\bigr{)}$ denotes the Kullback–Leibler (KL) divergence between the joint distribution $P_{X,W}$ and the product of the marginal distributions $P_{X}\otimes P_{W}$.
Theorem 1 of [1] lays the foundation of PAC Privacy analysis, which provides a quantitative upper bound on the adversary’s inference advantage in terms of $f$-divergence. Specifically, for any $f$-divergence $\mathtt{D}_{f}$, the adversary’s posterior advantage $\Delta_{f}\delta$ is bounded by the minimum $f$-divergence between the joint distribution $(X,\mathcal{M}(X))$ and the product of the marginal $P_{X}$ and any auxiliary output distribution $P_{W}$ that is independent of $X$:
$\Delta_{f}\delta\leq\inf_{P_{W}}\mathtt{D}_{f}\big{(}P{X,\mathcal{M}(X)}\;\|\;P_{X}\otimes P_{W}\big{)},$ (2)
where $P_{X,\mathcal{M}(X)}$ denotes the joint distribution of the data and mechanism output, respectively, and $P_{W}$ ranges over all distributions on the output space.
When the $f$-divergence is instantiated as the KL divergence and $P_{W}$ is chosen as the marginal output distribution $P_{\mathcal{M}(X)}$, this bound becomes the mutual information between $X$ and $\mathcal{M}(X)$:
$\Delta_{\mathrm{KL}}\delta\leq\mathtt{MI}(X;\mathcal{M}(X)),$
showing that the mutual information between the input and the mechanism output governs the worst-case posterior advantage.
For any selected $f$-divergence $\mathtt{D}_{f}$, the inequality (2) implies that a mechanism $\mathcal{M}:\mathcal{X}\rightarrow\mathcal{Y}$ satisfies $(\mathtt{R}_{f}^{\delta},\rho,\mathcal{D})$ Residual-PAC Privacy if
@@
Let $\mathtt{D}_{f}$ be the KL-divergence. $\mathcal{M}:\mathcal{X}\rightarrow\mathcal{Y}$ satisfies $(\mathtt{R}_{f}^{\delta},\rho,\mathcal{D})$ Residual-PAC Privacy if
@@
[…]
  • Proof.
[…]

### v2

@@ 4.1 Foundation of Residual-PAC Privacy
In this section, we develop general results to support concrete analyses under R-PAC Privacy framework. We begin by introducing key information-theoretic quantities, entropy and conditional entropy.
Entropy. The Shannon entropy of a discrete random variable $X$ on alphabet $\mathcal{X}$ is given by
$\mathcal{H}(X)=-\sum\nolimits_{x\in\mathcal{X}}P_{X}(x)\log P_{X}(x)$
while for continuous $X$, the differential entropy is
$h(X)=-\int\nolimits_{\mathcal{X}}f_{X}(x)\log f_{X}(x)\,dx.$
Conditional Entropy. For jointly distributed random variables $(X,W)$ where $X$ is on alphabet $\mathcal{X}$ and $W$ is on alphabet $\mathcal{W}$ (discrete case) or domain $\mathcal{W}$ (continuous case), the conditional entropy of $X$ given $W$ is defined as
$\mathcal{H}(X|W)=\sum\nolimits_{w\in\mathcal{W}}P_{W}(w)\mathcal{H}(X|W=w)$
in the discrete case and
$h(X|W)=\int\nolimits_{\mathcal{W}}f_{W}(w)h(X|W=w)\,dw$
in the continuous case, where $\mathcal{H}(X|W=w)=-\sum\nolimits_{x\in\mathcal{X}}P_{X|W}(x|w)\log P_{X|W}(x|w)$ and $h(X|W=w)=-\int_{\mathcal{X}}f_{X|W}(x|w)\log f_{X|W}(x|w)\,dx$.
For ease of exposition, we use $\mathcal{H}(X)$ to denote the entropy of $X$, either Shannon or differential depending on the context, and $\mathcal{H}(X|W)$ to denote the corresponding conditional entropy. When all entropies are finite, mutual information can equivalently be expressed as
$\mathtt{MI}(X;W)=\mathcal{H}(X)-\mathcal{H}(X|W).$
Consider any $f$-divergence $\mathtt{D}_{f}$, Theorem 1 of [46] shows that the posterior advantage $\Delta_{f}^{\delta}$ is bounded by the minimum $f$-divergence between the joint distribution of $(X,\mathcal{M}(X))$, denoted by $P_{X,\mathcal{M}(X)}$, and the product of the marginal distribution $P_{X}$ and any auxiliary output distribution $P_{W}$ independent of $X$:
$\Delta_{f}^{\delta}\leq\inf_{P_{W}}\mathtt{D}_{f}\big{(}P_{X,\mathcal{M}(X)}\;\|\;P_{X}\otimes P_{W}\big{)},$ (10)
where $P_{W}$ ranges over all distributions on the output space. where $P_{X,\mathcal{M}(X)}$ denotes the joint distribution of the data and mechanism output, and $P_{W}$ ranges over all distributions on the output space. When $\mathtt{D}_{f}$ is instantiated as $\mathtt{D}_{\mathtt{KL}}$ and $P_{W}=P_{\mathcal{M}(X)}$, we obtain (1).
Thus, for any $f$-divergence $\mathtt{D}_{f}$, inequality (10) implies that a mechanism $\mathcal{M}:\mathcal{X}\rightarrow\mathcal{Y}$ satisfies $(\mathtt{R}_{f}^{\delta},\rho,\mathcal{D})$ R-PAC Privacy if
@@
Let $R$ be a random variable distributed according to the reference distribution $\mathcal{R}$ over $\mathcal{X}$.
@@
Suppose that $\mathcal{H}(X)$ is finite and let $\mathtt{D}_{f}$ be the KL divergence. A mechanism $\mathcal{M}:\mathcal{X}\rightarrow\mathcal{Y}$ satisfies $(\mathtt{R}_{f}^{\delta},\rho,\mathcal{D})$ R-PAC Privacy if
@@
where $\mathtt{V}=\mathcal{H}(R)$ is the entropy of the reference distribution.